Diffusion Morning Report 05/03/06

Ted Berg tberg at snf.stanford.edu
Wed May 3 07:56:49 PDT 2006


Good  Morning All,
    Looking good this morning. Epi HCl crossed over. TEOS pump is bad 
and out for repair-approximately 1.5 weeks. Thermco2 back up after
removing burnt wipe. Everything else looks OK for now.

Diffusion 	  	  	  	  	 
  	Date: 	5/3/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Down 	Pump out for repair 	20.43/30 	  	 
Tylan 8-SiGe 	Up 	  	29.39/30 	  	 
Tylan 9-Poly 	Up 	  	2.73/30 	  	 
Tylan 10-Nitride 	Up 	  	24.51/30 	  	 
Tylan 11-BPSG-LTO 	up 	  	5.12/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Down 	Testing 	  	  	 
  	  	  	  	  	 
Epi 	yellow 	possible higher leak rate 	  	  	 
  	  	HCl Changed today 	  	  	 
Burn Box 	Up 	  	  	  	 
ZYGO 	up 	  	   	  	 
  	  	  	  	  	 
GAAS 22 	UP 	  	  	  	 
GAAS 23 	up 	  	  	  	 
  	  	  	  	  	 
SiH4 for furnaces 	  	A on line 575 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060503/400ad1c4/attachment.html>


More information about the diffusion mailing list